DocDB 8816 February 2013 Muon Operations Water Systems General: O2 sensors need to be calibrated. They may be faulty, and not giving reliable readings. Random system stops continue, and are not understood. EH1: Pool resistivity holding at 1.5 Mohm-cm, though a new sensor at a different point indicates 5.0. The system continues to operate well. Last filter change: 25jan2013. Last resin change: 5jul12, though that is uncertain. EH2: Resistivity about 1.1 Mohm-cm, down from 2.6 in early February. Early in February, the vacuum tank imploded and shut down the water system. Resistivity is slowing improving following the repair. Last filter change: 25jan2013. Last resin change: 5nov2012 EH3: Level sensors have failed, including backups. Until the level sensors are replaced (sometime in March), the level is being maintained "manually", with a short fill once per week, and the O2 sensors have been valved off to minimize water loss. Pool resistivity remains good, above 6 Mohm-cm, and the water system is otherwise working well. EH3 is operating with 1 and 0.2 micron filters. Last filter change: 25jan2013. Last resin change: 25nov2012. EH4: Working well. Frequent alarms do not indicate serious problems, and are distracting to shifters. DocDB 8752 is a draft proposal for replacing the level sensors. Muon PMTs We would like to replace some EH3 PMTs, though the improvements from replacing the FEE cards may have made this less important. This is still being studied. See DocDB 8805. Dead PMTs as of 20feb13: DVIF42 DVIB32 DVIB41 LCIG72 LVIA44 LVOEF2 FCIA61 FVIA43 FVIE41 FVOC43 FVOA43 FVOG31 FVOE32 See DocDB 8752 for a draft proposal for replacing PMTs. Muon Calibration No news RPC Performance There is a dead module in the middle of EH3. Around middle of week of 2/14, the trigger rate efficiency and noise rate of EH1 RPCs went down, in coincidence with total gas flow rate going up by 25%. The total flow rate went back down to 13% above normal after about 10 days. This is being investigated. RPC Gas System All three RPC gas systems are running quite smoothly. On 1/28/2013 there was a minor interruption for the EH2 gas system due to a blown fuse in the gas pressure control crate. Zhijian Zhang from DGUT replaced the fuse the next day; the system immediately returned to normal operation. C. Lu is remotely monitoring the gas chromatograph systems and digital bubbler systems from time to time. RPC HV No news Overall Muon Performance There are no firm indications of changes in performance of the muon system in its primary role as a veto. However, improvements in muon reconstruction efficiency in EH3 have been documented in DocDB 8805. These are attributed to the "touching" of FEE cables during the investigation of noise on Dec 15, followed by further improvements with the replacement of FEE boards on Jan 13 and 15.